a square is inscribed in the circle x square plus y square - 2 X + Y - 93 is equal to zero with the sides parallel to co-ordinate Axes. the vertices are

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
Centre of Circle = (1,-1/2)
       Radius = √94
       Slope of one diagonal of square = 1 (tan 45 )
       Slope of another diagonal = -1
       Vertex 1 = ( 1+√94 cos 45 , -1/2 + √94 sin 45 )
Vertex 2 = ( 1 - √94 cos45 , -1/2 - √94 sin45 )
Vertex 3 = ( 1 + √94 cos 135, -1/2 + √94 sin135 )
Vertex 4 = ( 1 - √94 cos 135, -1/2 - √94 sin135 ) 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)